OpenStudy (anonymous):

Write a radical in which the radicand consists of a coefficient and a variable. Use a coefficient that can be simplified best using the greatest perfect square method. Then, simplify the radical completely.

OpenStudy (anonymous):

\[\sqrt{50x^5}\]is my example

OpenStudy (anonymous):

then since \(50=25\times 2\) and \(x^5=x^4\times x\) you can rewrite this as \[\sqrt{50x^5}=\sqrt{25x^4\times 2x}=\sqrt{25x^4}\sqrt{2x}=5x\sqrt{2x}\]

OpenStudy (anonymous):

now i am sure you can make up your own example

OpenStudy (anonymous):

Thank you!!

OpenStudy (anonymous):

yw